if the perimeter of a square is 20 centimeters what is the length of the diagonal?
Found 2 solutions by fractalier, bmckenzie40:
Answer by fractalier(6550)

If the perimeter of a square is 20 cm, each side must be 5 cm. If you draw a square with a side of 5 cm, you will notice that drawing a diagonal makes a 45-45-90 right triangle whose sides are in the 1 to 1 to sqrt(2) ratio. Thus the diagonal must be 5 * sqrt(2) cm.

Answer by bmckenzie40(10)

If the perimeter of the square is 20 cm's then that means that each side of the square is 5 cm's.
For the diagonal it cuts the square into 2 equal triangles so you can apply pythagorean's theory.
a^2 + b^2 = c^2
a and b are the two sides of the square being 5 cm's and c is the diagonal.
5^2 + 5^2 = c^2
25 + 25 = c^2
50 = c^2
therefore the diagonal is the square root of 50 which is approx. 7